# Hovering Flight in Flapping Insects and Hummingbirds: A Natural Real-Time and Stable Extremum Seeking Feedback System

### v3

@@ IV Stability analysis of our natural hovering extremum seeking system and comparison with open-loop
Insect Open Loop ESC ($J=z^{2}$) ESC ($J=\dot{w}^{2}$)
Hawkmoth $-5.11e+05$ $-5.25e+05$ $-5.48e+05$
$-1.78e+04$ $-1.82e+04$ $-1.90e+04$
Cranefly $-1.50e+06$ $-1.60e+06$ $-1.68e+06$
$-9.28e+04$ $-9.60e+04$ $-1.04e+05$
Bumblebee $-1.20e+07$ $-1.26e+07$ $-1.27e+07$
$-1.02e+05$ $-1.07e+05$ $-1.08e+05$
Dragonfly $-4.14e+06$ $-4.39e+06$ $-4.52e+06$
$-1.41e+05$ $-1.60e+05$ $-1.82e+05$
Hoverfly $-7.81e+06$ $-8.04e+06$ $-8.11e+06$
$-1.20e+05$ $-1.24e+05$ $-1.25e+05$
Hummingbird $-3.33e+06$ $-4.01e+06$ $-4.38e+06$
$-2.00e+05$ $-2.40e+05$ $-2.62e+05$
@@ V Conclusion and future work
This paper provided a novel characterization (the first of its kind) of the hovering phenomenon in flapping insects and hummingbirds. By describing hovering in flapping insects and hummingbirds as an extremum seeking system (we call it a natural hovering ES system), we were able to start a line of research that may lead to solving all the puzzle pieces of hovering flight that existed for decades in previous literature (depicted in Figure 1). The proposed natural hovering ES system: (i) is very simple computationally; (ii) relies only on the natural oscillation of the wing combined with sensation-based feedback; (iii) real-time; (iv) stable; and (v) model-free (we used available models in literature to test the hypothesis but ES does not need modeling for implementation). We demonstrated the effectiveness of the proposed natural hovering ES via successful simulation trials demonstrating stabilization (even under delays and persistent noise in feedback measurements) in real-time. We also compared the proposed natural hovering ES to open-loop and PID approaches, demonstrating the clear advantages the proposed framework offers. Additionally, we provided stability analysis, which confirmed the observations provided in the simulations In the future, we aim at expanding the proposed research to include pitching dynamics, which is at the center of the stability-instability debate and corresponding claims in literature. We also seek experimental validation of natural hovering ES.

### v4

@@ IV Stability analysis of our natural hovering extremum seeking system and comparison with open-loop
Insect Open loop Optimal PID ESC ($J=z^{2}$) ESC ($J=\dot{w}^{2}$)
Hawkmoth $-5.11e{+05}$ $-4.81e{+05}$ $-5.25e{+05}$ $-5.48e{+05}$
$-1.78e{+04}$ $-1.66e{+04}$ $-1.82e{+04}$ $-1.90e{+04}$
Cranefly $-1.50e{+06}$ $-1.60e{+06}$ $-1.60e{+06}$ $-1.68e{+06}$
$-9.28e{+04}$ $-1.01e{+05}$ $-9.60e{+04}$ $-1.04e{+05}$
Bumblebee $-1.20e{+07}$ $-1.31e{+07}$ $-1.26e{+07}$ $-1.27e{+07}$
$-1.02e{+05}$ $-1.12e{+05}$ $-1.07e{+05}$ $-1.08e{+05}$
Dragonfly $-4.14e{+06}$ $-4.50e{+06}$ $-4.39e{+06}$ $-4.52e{+06}$
$-1.41e{+05}$ $-1.54e{+05}$ $-1.60e{+05}$ $-1.82e{+05}$
Hoverfly $-7.81e{+06}$ $-8.55e{+06}$ $-8.04e{+06}$ $-8.11e{+06}$
$-1.20e{+05}$ $-1.31e{+05}$ $-1.24e{+05}$ $-1.25e{+05}$
Hummingbird $-3.33e{+06}$ $-3.65e{+06}$ $-4.01e{+06}$ $-4.38e{+06}$
$-2.00e{+05}$ $-2.14e{+05}$ $-2.40e{+05}$ $-2.62e{+05}$
After considering the aforementioned smoothing approximation, we are ready to apply the VOC-based averaging in (14) and analyze stability via linearization and eigenvalues, which in turn characterizes the stability of the natural hovering ES system (11). The reader is referred to the code in [60] for the computations involved in (14), linearization about hovering equilibrium, and eigenvalues computation. However, here we present the end result, provided in Table 10.
Additionally, we made computations and included also in Table 10 the eigenvalues of stable open-loop hovering [46]; the reader is encouraged to recall relevant comments/discussions about open-loop stability provided in Subsection III.A. To obtain the optimal–PID eigenvalues, the averaged model was first derived exactly as in the open-loop case. Classical linear-control theory was then applied to incorporate the PID loop and evaluate the closed-loop eigenvalues, following the procedure in [59]. All computational steps are available in the accompanying code repository in [60]. It is important to mention that there is a difference in the nature of the stability between PID/ES on one hand and the open-loop case on the other. In the open-loop configuration, the equilibrium need not correspond to hovering ($w=0$). Stability can occur at a non-zero $w$, leading to steady ascent or descent, as discussed in Section III.B; only by selecting the exact trim (numerically demanding) does one obtain open-loop hovering. By contrast, both the PID loop and the ES controller always stabilize to $w\!=\!0$, so any disturbance is rejected and altitude is recovered. It can also be noted that in some instances the optimal PID closed-loop is less effective than the open-loop system—in the hawkmoth case, for example, as shown in Table 10 (but this should be put in the context that PID stabilization is hovering stabilization, which is not the case in the open-loop case. Overall, the eigenvalue results confirm the improvement in hovering stability by the natural hovering ES system using different objective functions compared to the open-loop case and optimal PID case. The results confirm that via extremum seeking control (ESC), hovering exhibits eigenvalues with more negative real components (i.e., shifts the eigenvalues further into the left-half complex plane), indicating stronger stability properties and more robustness.
@@ V Conclusion and future work
This paper provided a novel characterization (the first of its kind) of the hovering phenomenon in flapping insects and hummingbirds. By describing hovering in flapping insects and hummingbirds as an extremum seeking system (we call it a natural hovering ES system), we were able to start a line of research that may lead to solving all the puzzle pieces of hovering flight that existed for decades in previous literature. The proposed natural hovering ES system: (i) is very simple computationally; (ii) relies only on the natural oscillation of the wing combined with sensation-based feedback; (iii) real-time; (iv) stable; and (v) model-free (we used available models in literature to test the hypothesis but ES does not need modeling for implementation). We demonstrated the effectiveness of the proposed natural hovering ES via successful simulation trials demonstrating stabilization (even under delays and persistent noise in feedback measurements) in real-time. We also compared the proposed natural hovering ES to open-loop and PID approaches, demonstrating the clear advantages the proposed framework offers. Additionally, we provided stability analysis, which confirmed the observations provided in the simulations. In the future, we aim at expanding the proposed research to include pitching dynamics, which is at the center of the stability-instability debate and corresponding claims in literature. We also seek experimental validation of natural hovering ES.